6. Solve the following recurrence relations: (a) An+1 = 2 an , AO = 2 (b) n-1 An+1 =1+ ak , 0o = a1 = 1 ,n> 1 k=0
Answer #1

al antia on ho=2 a = (a) 22 A (21) age Go az 2 a ao an a la (agah hool ana pers 4 nzoAntalt 6 autza It he had anti- auta , nal an The recurrence relation can be written au - anya an-z=0 Here we use the charactero no 2 こ A0 = 1 = G(14YR) + 6(1-402* = 4+2 2121 = G ( 14YR) + CL-Y) Sorwing of Geltra 2 1 (1–0)( 4 1-5 =) 7 > VEGA VECE 사소 H